# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v2

@@ Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time white noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{3}{2},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
@@ 1.4 Main result
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s..
@@
‘a.s.’ represent almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represent the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
@@
For $\gamma\in(\frac{3}{2},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ where $\delta$ satisfies Theorem 1.1, Rough Burgers equation (1.1) has a global weak solution $u=v+X$ where $X$ satisfies (1.5) and $v$ is belongs to $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$ . Moreover, for $\gamma>\frac{5}{3}$, such solution is the global strong solution of (1.1).
@@
Theorem 1.1 and 1.2 is also valid for Rough DP equation. In fact, the nonlinear term $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$ in the DP equation does not introduce additional difficulties for the local well-posedness problem. For global results, we need to utilize the special structure of the DP equation to derive energy estimates as Proposition 3.3 and Lemma 3.4.
@@
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for $0<\delta<\frac{\gamma}{2}-\frac{5}{8}$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfied
@@
$\mathcal{L}u^{\sharp}=$ $\mathcal{L}u-\mathcal{L}X-\mathcal{L}u^{\prime}\prec\!\!\!\prec\mathcal{Q},$
@@
where $X(t)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}W^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Moreover, $u$ is a solution of (1.1).
@@
$\prec$ and $\prec\!\!\!\prec$ is referred to as the para-product and modified para-product, which will be introduced in Lemma 2.1 and (2.1). Let $v:=u^{\prime}\prec\!\!\!\prec\mathcal{Q}$, then $v$ satisfied the equation
$\partial_{t}v-\Lambda^{\gamma}v=^{\prime}regular\;term^{\prime}+u^{\prime}\prec\partial_{x}X$
This term that will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gains higher regularity (See Proposition 4.1), we will show it in Section 4.
@@
In fact, we can’t prove such result with only $H^{\alpha}$ space under our setting. Let $(u^{\prime},\mathcal{Q},u^{\sharp})\in(\mathcal{C}^{\alpha}\times H^{\beta}\times H^{\eta})$, then $\alpha+\beta=\eta$ is always valid. Since $\mathcal{Q}$ can gains $(\alpha-1+\gamma)$-order Sobolev regularity, then $u^{\sharp}$ gains $(2\alpha-1+\gamma)$-order Sobolev regularity at most by proof of Lemma B.3. To establish the $C^{\alpha}$ Hölder regularity of $u^{\prime}$, it is necessary for the remainder $u^{\sharp}$ to possess Sobolev regularity of order at least $H^{\alpha+\frac{1}{2}+}$, this is valid only when $\gamma>\frac{4}{3}$.
@@
We can extend the range to $\gamma\in(1,2]$, which means $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, since we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
@@ 2.1 Littlewood-Paley theory
$\mathcal{F}(\Delta_{j}u)(k)=\rho(2^{-l}k)u_{k},\quad l\geq 0\quad and\quad\mathcal{F}(\Delta_{-1}u)(k)=\mathcal{\chi}(k)u_{k}.$
@@
It can define $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{s}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and Sobolev space $H^{s}(\mathbb{T}^{d}):=B^{s}_{2,2}(\mathbb{T}^{d})$ such that
$B_{p,r}^{s}(\mathbb{T}^{d}):=\{u:(\sum_{l\geq-1}2^{lsr}\|\Delta_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r})^{\frac{1}{r}}<\infty\},$
also define the norm. Under this setting, we introduce para-product decomposition which give a product rules in $\mathcal{S}^{\prime}(\mathbb{T}^{d})$. Let $S_{j}u=\sum_{k\leq j-1}\Delta_{k}u$, then we have the estimate
@@ 2.3 Iteration argument
$<$ $(\frac{7}{4}+c)\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}<\frac{15}{16}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}$
Then we get the limit of $(x_{n},y_{n})$ is a unique fixed point of mapping (2.7) in Banach space $\mathcal{E}$, which prove the lemma. ∎
@@ 3.1 Bi-linear estimate
Let $\gamma\in(\frac{4}{3},2]$, $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\delta$ satisfy $0<2\delta\leq\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have the following bounds
@@ 3.2 Well-posedness
Let $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta\leq\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{\bar{T}}W^{\alpha}}$ for some $\bar{T}>0$ , then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfying $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
@@
$(\int_{\mathbb{T}^{d}}\langle\xi\rangle^{2s}(\sup_{0\leq t^{\prime}\leq t}|\hat{v}(t^{\prime},\xi)|)^{2}\mathrm{d}\xi)^{\frac{1}{2}}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(\frac{1}{4},\frac{\gamma}{2}-\frac{1}{2})$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. Then for any $t\in[0,T]$, we have the following energy estimate
$\|v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})t}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})(t-t^{\prime})}\mathrm{d}t^{\prime}.$ (3.7)
@@
$\|v(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+2\int_{0}^{t}\langle v(t^{\prime}),2\partial_{x}(v(t^{\prime})\tilde{X}(t^{\prime}))+\partial_{x}(\tilde{X}^{2}(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality, Young’s inequality, Lemma 2.1 and Lemma 2.7, there exist $\alpha\in(\frac{1}{4},\frac{1}{2})$ such that for $\frac{\gamma}{2}-\frac{1}{2}>\alpha>1-\frac{\gamma}{2}$, we have estimate
@@
$\|v(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}})\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}$ (3.8)
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $s<\gamma-\frac{3}{2}$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
@@
$+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime},$
for some small $\nu<1$ and any $0<s<\gamma-\frac{3}{2}$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
@@
For the first term in the right side of (3.11), by Parsavel formula and Hölder inequality, we have
@@
By the estimate in [KPV93, GO14] and the interpolation, for $0<s+1-\frac{\gamma}{2}<1$ and $\alpha=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have

### v3

@@ Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time white noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{5}{3},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
@@ 1.4 Main result
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ satisfies (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s..
@@
‘a.s.’ represents almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represents the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
@@
Let $\gamma\in(\frac{5}{3},2]$, $u_{0}\in H^{\sigma}$ for $\sigma>0$ and $\int_{\mathbb{T}}u(x)\mathrm{d}x=0$, then Rough Burgers equation (1.1) has a global mild solution $u=v+X$ where $X$ satisfies (1.5) and $v$ belongs to $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$.
@@
Theorem 1.1 and 1.2 are also valid for Rough DP equation. In fact, the nonlinear term $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$ in the DP equation does not introduce additional difficulties for the local well-posedness problem. For global results, we need to utilize the special structure of the DP equation to derive energy estimates as Proposition 3.3 and Lemma 3.4.
@@
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for $0<\delta<\frac{\gamma}{2}-\frac{5}{8}$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfying
@@
$\mathcal{L}u^{\sharp}=$ $\mathcal{L}u-\mathcal{L}X-\mathcal{L}u^{\prime}\prec\!\!\!\prec\mathcal{Q},\quad u^{\sharp}(0)=u_{0}$
@@
where $X(t)$ satisfies (1.5) with $X(t,x)\in C_{T^{*}}W^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Moreover, $u$ is a solution of (1.1).
@@
$\prec$ and $\prec\!\!\!\prec$ are referred to as the para-product and modified para-product, which will be introduced in Lemma 2.1 and (2.2). Let $v:=u^{\prime}\prec\!\!\!\prec\mathcal{Q}$, then $v$ satisfies the equation
$\partial_{t}v-\Lambda^{\gamma}v=^{\prime}regular\;term^{\prime}+u^{\prime}\prec\partial_{x}X.$
This term will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gain higher regularity (See Proposition 4.1). We will show it in Section 4.
@@
In fact, we can’t prove such result with only $H^{\alpha}$ space under our setting. Let $(u^{\prime},\mathcal{Q},u^{\sharp})\in(\mathcal{C}^{\alpha}\times H^{\beta}\times H^{\eta})$, then $\alpha+\beta=\eta$ is always valid. Since $\mathcal{Q}$ can gain $(\alpha-1+\gamma)$-order Sobolev regularity, $u^{\sharp}$ gains $(2\alpha-1+\gamma)$-order Sobolev regularity at most by proof of Lemma B.3. To establish the $C^{\alpha}$ Hölder regularity of $u^{\prime}$, it is necessary for the remainder $u^{\sharp}$ to possess Sobolev regularity of order at least $H^{\alpha+\frac{1}{2}+}$, this is valid only when $\gamma>\frac{4}{3}$.
@@
We can extend the range to $\gamma\in(1,2]$, which means $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
@@ 2.1 Littlewood-Paley theory
$\mathcal{F}(\Delta_{l}u)(k)=\rho(2^{-l}k)u_{k},\quad l\geq 0\quad and\quad\mathcal{F}(\Delta_{-1}u)(k)=\mathcal{\chi}(k)u_{k}.$
@@
It can define nonhomogeneous Besov space $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{s}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and nonhomogeneous Sobolev space $H^{s}(\mathbb{T}^{d}):=B^{s}_{2,2}(\mathbb{T}^{d})$ such that
$B_{p,r}^{s}(\mathbb{T}^{d}):=\left\{u\in S^{\prime}:\left(\sum_{l\geq-1}2^{lsr}\|\Delta_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r}\right)^{\frac{1}{r}}<\infty\right\},$
also define the norm. Also define the Fourier multiplier
$\mathcal{F}(\dot{\Delta}_{j}u)(k)=\rho(2^{-l}k)u_{k},\quad l\in\mathbb{Z},$
where $\rho$ is a smooth cut-off function supported in an annulus. Then we introduce the homogeneous Besov space
$\dot{B}_{p,r}^{s}(\mathbb{T}^{d}):=\left\{u\in\mathcal{S}_{h}^{\prime}:\left(\sum_{l\in\mathbb{Z}}2^{lsr}\|\dot{\Delta}_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r}\right)^{1/r}<\infty\right\},$
where $\mathcal{S}_{h}^{\prime}$ denotes the space of tempered distributions $u$ such that for any $\theta\in\mathcal{D}(\mathbb{T}^{d})$ with value $1$ near 0,
$\lim_{\lambda\to\infty}\|\theta(\lambda D)u\|_{L^{\infty}}=0.$ (2.1)
It can be verified that $\dot{B}_{2,2}^{s}(\mathbb{T}^{d})$ coincides with the homogeneous Sobolev space
$\dot{H}^{s}(\mathbb{T}^{d})=\left\{u\in\mathcal{S}_{h}^{\prime}:\left(\sum_{k\neq 0}|k|^{2s}|\mathcal{F}(u)(k)|^{2}\right)^{1/2}<\infty\right\}.$
It is straightforward to verify that condition (2.1) implies $\int_{\mathbb{T}}u(x)\,dx=\widehat{u}(0)=0$. Consequently, a function $f$ belongs to $\dot{H}^{s}$ if and only if $f\in H^{s}$ and $\int_{\mathbb{T}}f(x)\,dx=0$. Under this setting, we introduce para-product decomposition which give a product rules in $\mathcal{S}^{\prime}(\mathbb{T}^{d})$. Let $S_{j}u=\sum_{k\leq j-1}\Delta_{k}u$, then we have the estimate
@@
The above estimate also holds if we replace $\dot{\Delta}_{j}$ with $\Delta_{j}$ and $\dot{S}_{j}$ with $S_{j}$.
@@ 2.3 Iteration argument
$<$ $(\frac{7}{8}+\frac{13c}{8})\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}<\frac{15}{16}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}.$
Then we get the limit of $(x_{n},y_{n})$ is a unique fixed point of mapping (2.8) in Banach space $\mathcal{E}$, which proves the lemma. ∎
@@ 3.1 Bi-linear estimate
Let $\gamma\in(\frac{4}{3},2]$, $\alpha>\frac{1}{6}$, $\delta$ satisfy $0<2\delta\leq\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have the following bounds
@@ 3.2 Well-posedness
Let $\alpha>\frac{1}{6}$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta\leq\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{\bar{T}}W^{\alpha}}$ for some $\bar{T}>0$ , then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfying $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
@@
$(\sum_{k\neq 0}\langle k\rangle^{2s}(\sup_{0\leq t^{\prime}\leq t}|\hat{v}(t^{\prime},\xi)|)^{2})^{\frac{1}{2}}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
@@
Next, we build the well-posedness in $L^{4}_{T}\dot{H}^{s+\frac{\gamma}{4}}$, for $0\leq s<\gamma-\frac{3}{2}$. The key observation is that for $s_{1}\leq s_{2}$, we have
$\|\partial_{x}f\|_{\dot{H}^{s_{1}}}\leq\|\partial_{x}f\|_{\dot{H}^{s_{2}}}.$ (3.7)
@@
Let $u_{0}\in\dot{H}^{s}$ for $s\in[0,\gamma-\frac{3}{2})$ with $\gamma\in(\frac{8}{5},2]$, then there exists a constant $C$ such that if the following conditions
$CT_{0}^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}<\frac{1}{4}\quad for\;0<T_{0}\leq\bar{T}\quad and\quad\alpha\geq\max(\frac{3}{2}-\frac{3\gamma}{4},s+1-\frac{\gamma}{2}),$ (3.8)
hold, then there exists a positive time $T$ such that (3.3) has a unique solution in $L_{T}^{4}([0,T];\dot{H}^{s+\frac{\gamma}{4}})$. Moreover let $T_{u_{0}}$ denote the maximal time of existence of such a solution, then
  • For the same constant $C$,
    $\|u_{0}\|_{\dot{H}^{s}}\leq\frac{1}{16C}\Longrightarrow T_{u_{0}}=T_{0}.$
  • If $T_{u_{0}}\leq T_{0}$, then
    $\int_{0}^{T_{u_{0}}}\|u(t)\|_{\dot{H}^{s+\frac{\gamma}{4}}}^{4}dt=\infty.$ (3.9)
Proof.
We firstly give the local result for small initial data $u_{0}$. Noting that $B(f,g)$ is the solution of
$\partial_{t}B(f,g)-\Lambda^{\gamma}B(f,g)=\partial_{x}(fg).$
Let $\gamma>\frac{8}{5}$ and $0\leq s<\gamma-\frac{3}{2}$, by Lemma 3.3, (3.7) and Lemma 2.1, we have
$\|P(t)u_{0}+B(\tilde{X},\tilde{X})\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $\|u_{0}\|_{\dot{H}^{s}}+CT^{\frac{1}{2}}\|\tilde{X}\|_{C_{T}W^{\alpha}}^{2},$ (3.10)
$2\|B(v,\tilde{X})\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $CT^{\frac{1}{4}}\|\tilde{X}\|_{C_{T}W^{\alpha}}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}},$ (3.11)
$\|B(v,v)\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}.$ (3.12)
Let $r=\frac{1}{8C}$, we know that if $\|u_{0}\|_{\dot{H}^{s}}\leq\frac{1}{16C}$ and
$CT_{0}^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}<\frac{1}{4},\quad$
by Lemma 2.8, there exists a unique solution of (3.3) in the ball with center 0 and radius $\frac{1}{4C}$ in the space $L^{4}([0,T_{0}];\dot{H}^{s+\frac{\gamma}{4}})$. Now we consider the case of large initial data $u_{0}$ in $\dot{H}^{s}$. We split $u_{0}$ into small part in $\dot{H}^{s}$ and a large part with compactly supported Fourier transform. For that we fix some positive real number $\rho_{u_{0}}$ such that
$(\sum_{|k|\geq\rho_{u_{0}}}|k|^{2s}|\widehat{u}_{0}(k)|^{2})^{\frac{1}{2}}\leq\frac{1}{32C}.$
Then by $\|P(t)u_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq\|u_{0}\|_{\dot{H}^{s}}$ and $u_{0}^{\flat}=\mathcal{F}^{-1}(1_{B(0,\rho_{u_{0}})}\widehat{u}_{0})$, we get
$\|P(t)u_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $\frac{1}{32C}+\|P(t)u^{\flat}_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}$
$\leq$ $\frac{1}{32C}+\rho_{u_{0}}^{\frac{\gamma}{4}}\|P(t)u^{\flat}_{0}\|_{L_{T}^{4}H^{s}}$
$\leq$ $\frac{1}{32C}+(\rho_{u_{0}}^{\gamma}T)^{\frac{1}{4}}\|u_{0}\|_{\dot{H}^{s}}.$
Thus if
$T\leq\left(\frac{1}{\rho_{u_{0}}^{\frac{\gamma}{4}}32C\|u_{0}\|_{\dot{H}^{s}}}\right)^{4}\wedge T_{0},$ (3.13)
then we have the existence of a unique solution in the ball with center 0 and radius $\frac{1}{4C}$ in the space $L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}$. Noting that if $v$ is a solution in $L^{4}([0,T];\dot{H}^{s+\frac{\gamma}{4}})$, by Lemma 3.3, we have $v\in\mathcal{C}([0,T];\dot{H}^{s})\cap L^{2}([0,T];\dot{H}^{s+\frac{\gamma}{2}})$
Finally, we prove the blow-up criterion. Assume that we have a solution $v$ of (3.3) on the time interval $[0,T[$ such that
$\int_{0}^{T}\|v(t)\|_{\dot{H}^{s+\frac{\gamma}{4}}}^{4}dt<\infty,\quad T<T_{0}<\bar{T}.$
We claim that the lifespan $T_{u_{0}}$ of $v$ is greater than $T$. Indeed, thanks to Lemma 3.3, we have
$\sum_{k\neq 0}|k|^{2s}(\sup_{t\in[0,T[}|\widehat{v}(t,k)|)^{2}\leq$ $\|v_{0}\|_{H^{s}}+\|\partial_{x}(v^{2}+2v\tilde{X}+\tilde{X}^{2})\|_{\dot{H}^{s-\frac{\gamma}{2}}}$
$\leq$ $\|v_{0}\|_{H^{s}}+C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}+CT^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}+CT^{\frac{1}{2}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}$
$\leq$ $\|v_{0}\|_{H^{s}}+C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}+\frac{1}{4}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}+\frac{1}{16C}<\infty.$
Thus, there exists a positive number $\rho$ exists such that
$\forall t\in[0,T[,\quad\sum_{|k|\geq\rho}|k|^{2s}(\sup_{t\in[0,T[}|\widehat{v}(t,k)|)^{2}\leq\frac{1}{32C}.$
Noting that the choise of $\rho$ is independent of $t$, by condition of (3.13), we can get $T_{0}\geq T_{u_{0}}>T$ which finish our proof. ∎
Proposition 3.3.
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ and $0<T\leq\bar{T}$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in L^{2}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. Then for any $t\in[0,T]$, we have the following energy estimate
$\|v\|_{L^{2}}^{2}+2(1-\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})t}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})(t-t^{\prime})}\mathrm{d}t^{\prime}.$ (3.14)
@@
$\|v(t)\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+2\int_{0}^{t}\langle v(t^{\prime}),2\partial_{x}(v(t^{\prime})\tilde{X}(t^{\prime}))+\partial_{x}(\tilde{X}^{2}(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality, Young’s inequality, Lemma 2.1 and Lemma 2.7, there exists $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ such that
@@
$\|v(t)\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}$ (3.15)
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $0<s<\gamma-\frac{3}{2}$ and $0<T\leq\bar{T}$. Let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{s}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
@@
$+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime}$
for some small $\nu<1$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
@@
For the first term in the right side of (3.18), by Parsavel formula and Hölder’s inequality, we have
@@
By the estimate in [KPV93, GO14] and the interpolation, for $0<s+1-\frac{\gamma}{2}<1$ and $\beta=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have